# Hyper-Compression: Model Compression via Hyperfunction

### v3

@@ 5 Experiment and Analysis
LLaMA2-7B et al. (2023) FP16 12.50GB
Sheared-LLaMA-1.3B Xia et al. (2024) FP32 5.01GB
TinyLLaMA Zhang et al. (2024) FP32 4.09GB
@@
Table 3: The information of our used models as the testbed of different model compression methods.
This section illustrates the efficacy of our novel compression methodology on three widely utilized representative models: LLaMA2 series et al. (2023), UNet Ronneberger et al. (2015), and MobileNetV3 Howard et al. (2019), which correspond to large, middle, and small models, respectively. Table 3 shows the information of our used models as the testbed of different model compression methods. Favorably, our compression technique does not require post-hoc retraining, even at high compression ratios. For example, we achieve up to a 7.87×7.87$\times$7.87 × reduction in the model size of UNet within $1\%$ performance drop. This is a significant improvement over traditional methods that often require extensive retraining to restore efficacy. This capability sets a new benchmark in model compression. Because our proposed hyper-compression method requires no post-hoc training, we do not compare it with distillation-based methods that demands a great amount of data to train the student model.
Preferable compression ratio. As shown in Table 4, we apply the same pruning method used in Sheared-LLaMA-1.3B, TinyLLaMA, and LiteLLaMA, combined with our compression technique. We evaluate our compressed models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. The average test results of eight downstream tasks are shown in Table 4, where “File Size” indicates the size of the file that stores the model. Notably, our method can compress LLaMA2-7B by a factor of 2.60×2.60$\times$2.60 × while maintaining the average score decrease within 1%, which achieves the best balance, while Oother models either achieve sup-optimal compression rates or bear a large performance drop. For example, Sheared-LLaMA-1.3B achieves 2.50×2.50$\times$2.50 × with $15.57\%$ performance decreases, while TinyLLaMA achieves 3.60×3.60$\times$3.60 × with $14.33\%$ performance loss on average.
Table 4: Comparison between Our Methods and other compressed Models of LLaMA2-7B on Individual Evaluation Dataset. As a reference, the compression ratios of INT4 quantization in LLaMA-7B is 4×4$\times$4 ×
@@
LLaMA2-7B et al. (2023) 12.50 94.00 68.98 76.30 52.39
Sheared-LLaMA-1.3B Xia et al. (2024) 5.01 (2.50$\times$) 87.30 58.09 60.98 34.04
TinyLLaMA Zhang et al. (2024) 4.09 (3.06$\times$) 89.30 59.43 61.66 37.12
LiteLLaMA 11footnotemark: 1 0.86 (14.53$\times$) 75.10 52.64 47.85 24.32
OPT-1.3B 2.63 (4.75$\times$) 84.30 59.60 57.00 29.70
Pythia-1.4B 2.93 (4.27$\times$) 86.40 57.40 60.70 31.20
OPT-2.7B 5.30 (2.36$\times$) 85.80 60.80 60.80 34.00
Pythia-2.8B 5.68 (2.20$\times$) 88.30 59.70 64.40 36.40
INCITE-Base-3B 5.69 (2.20$\times$) 90.70 63.50 67.70 40.20
Open-LLaMA–3B-v1 6.85 (1.82$\times$) 91.30 61.50 67.60 39.60
Open-LLaMA–3B-v2 6.85 (1.82$\times$) 91.80 63.50 66.50 39.00
LLaMA2-7B + HF 4.80 (2.60$\times$) 93.70 69.77 75.59 53.24
Sheared-LLaMA-1.3B + HF 0.98 (12.76$\times$) 87.90 58.88 60.48 32.85
TinyLLaMA + HF 0.78 (16.03 $\times$) 89.50 58.96 61.11 36.43
LiteLLaMA + HF 0.39 (32.05$\times$) 73.00 54.22 44.78 23.89
@@
LLaMA2-7B et al. (2023) 78.94 81.90 28.67 45.86 65.88
Sheared-LLaMA-1.3B Xia et al. (2024) 61.02 65.54 9.89 25.59 50.31
TinyLLaMA Zhang et al. (2024) 62.48 62.91 12.52 26.79 51.53
LiteLLaMA 38.41 57.09 1.77 26.11 40.41
@@
LLaMA2-7B + HF 77.17 80.92 25.65 43.04 64.89
Sheared-LLaMA-1.3B + HF 60.56 63.76 8.98 24.68 49.76
TinyLLaMA + HF 61.92 58.41 11.58 27.28 50.65
LiteLLaMA + HF 37.56 56.57 1.16 26.61 39.72
Model File Size (MB) Dice (%)
UNet Ronneberger et al. (2015) 51.10 99.86
Pruning 20.30 (2.53$\times$) 96.34
HF 6.49 (7.87$\times$) 99.71
Pruning + HF 2.88 (17.74$\times$) 96.45
Model File Size (MB) Top-1 Accuracy (%)
MobileNetV3 Howard et al. (2019) 5.95 74.41
Pruning 2.22 (2.68$\times$) 69.32
HF 1.65 (3.61$\times$) 73.93
Pruning + HF 0.47 (12.66$\times$) 68.47
Table 5: The compression effectiveness of our method on UNet and MobileNetV3. As a reference, the compression ratios of INT4 quantization in UNet, and MobileNetV3 are 8×8$\times$8 ×, and 8×8$\times$8 ×, respectively.
Moreover, perplexity (PPL) is a vital metric for evaluating the overall performance of large language models (LLMs). We conduct tests on the publicly available dataset wikitext [^2] For more details, please visit the Hugging Face website of INCITE-Base: https://huggingface.co/togethercomputer/RedPajama-INCITE-Base-3B-v1.. As shown in Table 6, the PPL of LLaMA2-7B+HF increases by only 0.35 compared to the original LLaMA2-7B model, whereas the PPL increases caused by other three LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA) are 2.66, 2.24, and 26.35, respectively. We also compare the perplexity with four quantization methods: W8A8 SmoothQuant (int8, $\alpha$=0.85), Asymmetric Quant (int8), GPTQ (int4), and Asymmetric Quant (int4). Among them, both W8A8 SmoothQuant (int8, $\alpha$=0.85) and GPTQ (int4) require calibration sets for calibration during the quantization process, whereas our method does not involve calibration. Without re-training, our model is already better than the int8 quantization but inferior to the int4 quantization with retraining. This demonstrates that our method is competitive in real-world applications. Figure 10 compares the text outputs generated by eight different LLM models for a given prompt.
@@
LLaMA2-7B + W8A8 SmoothQuant Xiao et al. (2023) 2.00$\times$ 5.52
LLaMA2-7B + Asymmetric Quant (int8) 2.00$\times$ 5.65
LLaMA2-7B + GPTQ Frantar et al. (2022) (int4) 3.46$\times$ 5.69
LLaMA2-7B + Asymmetric Quant (int4) 4.00$\times$ 26160.34
LLaMA2-7B + HF 2.60$\times$ 5.82
Sheared-LLaMA-1.3B 2.50$\times$ 8.13
Sheared-LLaMA-1.3B + HF 12.76$\times$ 8.37
TinyLLaMA 3.06$\times$ 7.71
TinyLLaMA + HF 16.03$\times$ 7.95
LiteLLaMA 14.53$\times$ 31.82
LiteLLaMA + HF 32.05$\times$ 37.85
Table 6: The comparison of Perplexity (PPL) on the dataset wikitext-2-raw-v1

### v4

@@ 5 Experiment and Analysis
TABLE III: The information of our used models as the testbed of different model compression methods.
@@
LLaMA2-7B [17] FP16 12.50GB
Sheared-LLaMA-1.3B [51] FP32 5.01GB
TinyLLaMA [56] FP32 4.09GB
@@
LLaMA-3.2-1B-Instruct  [^4] For more details, please visit the Hugging Face website of LLaMA-3.2-1B-Instruct: https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct. FP16 2.47GB
Qwen1.5-7B-Chat  [^5] For more details, please visit the Hugging Face website of Qwen1.5-7B-Chat: https://huggingface.co/Qwen/Qwen1.5-7B-Chat. FP16 15.45GB
DeepSeek-R1-Distill-Qwen-14B  [^6] For more details, please visit the Hugging Face website of DeepSeek-R1-Distill-Qwen-14B and 32B: https://huggingface.co/deepseek-ai. FP16 29.54GB
DeepSeek-R1-Distill-Qwen-32B 66footnotemark: 6 FP16 65.54GB
Qwen2-72B-Instruct [48] FP16 145.51GB
TABLE IV: Detailed hyperparameter settings in Hyper-Compression for all the models.
Model Max class $U$ $l$
UNet 3 [225] [0.1]
MobileNetV3 4 [1600] [0.8]
LLaMA2-7B 3 [1600] [0.1]
Sheared-LLaMA-1.3B 3 [1600] [0.1]
TinyLLaMA 3 [1600] [0.1]
LiteLLaMA 3 [1600] [0.1]
DeepSeek-R1-Distill-QWen-14B 3 [1681] [0.1]
DeepSeek-R1-Distill-QWen-32B 3 [1681] [0.1]
DeepSeek-R1-Distill-QWen-72B 3 [900, 1296, 1521, 1681] [0.1]
LLaMA-3.2-1B-Instruct 3 [1681] [0.1]
Qwen1.5-7B-Chat 3 [900, 1296, 1521, 1681] [0.05, 0.1, 0.15]
This section illustrates the efficacy of our novel compression methodology on three widely utilized representative models: LLaMA [17] series, Qwen [1] series, UNet [42], ResNet series [25], and MobileNetV3 [26], which correspond to large, middle, and small models, respectively. Table III shows the information of our used models as the testbed of different model compression methods. Favorably, our compression technique does not require post-hoc retraining, even at high compression ratios. For example, we achieve up to a $7.87\times$ reduction in the model size of UNet within $1\%$ performance drop. This is a significant improvement over traditional methods that often require extensive retraining to restore efficacy. This capability sets a new benchmark in model compression. Because our proposed Hyper-Compression method requires no post-hoc training, we do not compare it with distillation-based methods that demands a great amount of data to train the student model.
[…]
@@
LLaMA2-7B [17] 12.50 94.00 68.98 76.30 52.39
DeepSeek-R1-Distill-QWen-14B 66footnotemark: 6 29.54 95.50 72.85 78.11 58.62
DeepSeek-R1-Distill-QWen-32B 66footnotemark: 6 65.54 95.90 75.69 81.06 63.74
Qwen2-72B-Instruct [48] 145.51 97.10 77.82 85.27 68.60
OPT-1.3B 2.63 84.30 59.60 57.00 29.70
Pythia-1.4B 2.93 86.40 57.40 60.70 31.20
OPT-2.7B 5.30 85.80 60.80 60.80 34.00
Pythia-2.8B 5.68 88.30 59.70 64.40 36.40
INCITE-Base-3B 5.69 90.70 63.50 67.70 40.20
Open-LLaMA–3B-v1 6.85 91.30 61.50 67.60 39.60
Open-LLaMA–3B-v2 6.85 91.80 63.50 66.50 39.00
[…]
TinyLLaMA + HC 0.78 (16.03 $\times$) 89.50 58.96 61.11 36.43
[…]
@@
LLaMA2-7B [17] 78.94 81.90 28.67 45.86 65.88
DeepSeek-R1-Distill-QWen-14B 66footnotemark: 6 81.05 88.69 22.22 74.62 71.46
DeepSeek-R1-Distill-QWen-32B 66footnotemark: 6 83.27 90.52 26.37 80.90 74.68
Qwen2-72B-Instruct [48] 86.95 89.88 36.59 83.91 78.27
@@
Sheared-LLaMA-1.3B [51] 61.02 65.54 9.89 25.59 50.31
TinyLLaMA [56] 62.48 62.91 12.52 26.79 51.53
LiteLLaMA 38.41 57.09 1.77 26.11 40.41
LLaMA2-7B + HC 77.17 80.92 25.65 43.04 64.89
Sheared-LLaMA-1.3B + HC 60.56 63.76 8.98 24.68 49.76
TinyLLaMA + HC 61.92 58.41 11.58 27.28 50.65
LiteLLaMA + HC 37.56 56.57 1.16 26.61 39.72
DeepSeek-R1-Distill-QWen-14B + HC 80.36 88.44 21.39 73.47 70.50
DeepSeek-R1-Distill-QWen-32B + HC 82.94 90.34 25.18 80.43 74.31
Qwen2-72B-Instruct + HC 86.85 89.85 36.29 83.73 78.31
TABLE VI: The comparison of Perplexity (PPL) on the dataset wikitext-2-raw-v1
@@
[…]
Hardware Configuration and Detailed Compression Parameters. All experiments are conducted on a server equipped with NVIDIA RTX L40, NVIDIA RTX A40, NVIDIA RTX 4060 and NVIDIA RTX 6000 Pro GPUs. Given the extensive scope of our experiments, listing all hyperparameters in the main text would be exhaustive. Therefore, we have compiled a detailed Hyperparameter Table IV, which covers the specific settings for all reported results.
Preferable compression ratio. As shown in Table V, we apply the same pruning method used in Sheared-LLaMA-1.3B, TinyLLaMA, and LiteLLaMA, combined with our compression technique. We evaluate our compressed models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. The average test results of eight downstream tasks are shown in Table V, where “File Size” indicates the size of the file that stores the model. Notably, our method can compress LLaMA2-7B by a factor of $2.60\times$ while maintaining the average score decrease within 1%, which achieves the best balance, while other models either achieve sup-optimal compression rates or bear a large performance drop. For example, Sheared-LLaMA-1.3B achieves $2.50\times$ with $15.57\%$ performance decreases, while TinyLLaMA achieves $3.60\times$ with $14.33\%$ performance loss on average. Furthermore, we test the compressed model and the original one on more downstream benchmarks. The results are shown in Table VII. We expand the evaluation to 9 new tasks: GSM8K[11] for multi-step mathematical reasoning; PIQA[5] for physical commonsense understanding; and the SuperGLUE suite (CB[13], COPA[41], MultiRC[30], ReCoRD[57], RTE[44], WiC[39], WSC[32]) covering linguistic inference, coreference resolution, and reading comprehension. These tasks collectively assess critical dimensions of language model performance, including logical reasoning, world knowledge, and contextual understanding, which can provide a more rigorous validation of model robustness. To the best of our knowledge, few studies in the domain of model compression tested as many as 16 downstream tasks as our work does. For instance, Sheared-LLaMA [51] evaluated 11 downstream tasks, and AWQ [33] included 11 tasks in its comparative experiments; meanwhile, LLaMA [17] and Qwen2.5 [1] were tested on only 8 and 5 tasks, respectively.
TABLE VII: Performance comparison of Qwen1.5-7B-Chat and its hyper-compressed variant across 16 benchmark tasks
Model Ratio SciQ WinoGrande ARC-E ARC-C HellaSwag BoolQ MMLU GSM8K
Qwen1.5-7B-Chat - 92.60 65.43 68.48 52.56 78.65 85.44 60.48 54.06
+ Hyper-Compression 2.67 91.30 65.19 67.13 52.73 77.72 85.17 59.24 51.18
PIQA CB COPA MultiRC ReCoRD RTE WiC WSC Average Drop
75.19 60.71 84.00 26.26 80.18 83.03 67.40 55.77 65.23 -
74.65 60.71 89.00 23.54 79.87 80.59 66.77 59.62 64.75 0.48
TABLE VIII: The compression effectiveness of our method on ResNet18, ResNet50, UNet and MobileNetV3. Moreover, we also compare our method with other data-free methods on ResNet18 and ResNet50.
Model Ratio Dice (%)
UNet [42] - 99.86
[…]
Model Ratio Top-1 Acc (%)
MobileNetV3 [26] - 74.41
[…]
Model Ratio Top-1 Acc (%)
ResNet18 - 71.47
[…]
ResNet50 - 77.74
[…]
Moreover, perplexity (PPL) is a vital metric for evaluating the overall performance of large language models (LLMs). We conduct tests on the publicly available dataset wikitext [^2] For more details, please visit the Hugging Face website of INCITE-Base: https://huggingface.co/togethercomputer/RedPajama-INCITE-Base-3B-v1.. As shown in Table VI, the PPL of LLaMA2-7B+HC increases by only 0.35 compared to the original LLaMA2-7B model, whereas the PPL increases caused by other three LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA) are 2.66, 2.24, and 26.35, respectively. We also compare the perplexity with four quantization methods: W8A8 SmoothQuant (int8, $\alpha$=0.85), Asymmetric Quant (int8), GPTQ (int4), and Asymmetric Quant (int4). Among them, both W8A8 SmoothQuant (int8, $\alpha$=0.85) and GPTQ (int4) require calibration sets for calibration during the quantization process, whereas our method does not involve calibration. Without re-training, our model is already better than the int8 quantization but inferior to the int4 quantization with retraining. This demonstrates that our method is competitive in real-world applications. Figure 11 compares the text outputs generated by eight different LLM models for a given prompt.
We also incorporate comparative experiments with four prominent quantization methods: AWQ [33], GPTQ [20], SmoothQuant [52], and SpinQuant [35]. All four methods are Post-Training Quantization (PTQ) algorithms that even require calibration datasets for parameter optimization. We conduct these evaluations on the Llama-3.2-1B-Instruct and Qwen1.5-7B-Chat models.
The evaluation results are reported in Tables IX and X. We evaluate all the models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. Notably, our method compresses LLaMA-3.2-1B-Instruct and Qwen1.5-7B-Chat by factors of $1.62\times$ and $2.67\times$, respectively, while limiting the average score decrease to within 1.24% and 0.94%. This achieves the optimal balance between efficiency and accuracy, whereas other models either yield sub-optimal compression rates or suffer from significant performance degradation. For instance, although GPTQ-INT8 restricts the average score drop in LLaMA-3.2-1B-Instruct within only 0.06%, it achieves a limited compression factor of $1.69\times$. Similarly, while SmoothQuant-INT4 reaches a compression factor of $1.59\times$, it results in a substantial average performance drop of 5.14%. Considering that the INT4 quantization is the widely-recognized frontier, we reasonably conclude that our method is competitive. It should be noted that, owing to the extensive experimental scope, all comparative methods were evaluated using officially released open-source implementations from HuggingFace. This approach eliminates potential discrepancies arising from manual reimplementation, thereby ensuring the fairness of model performance assessment.
TABLE IX: Comparison between our methods and other compressed models of LLaMA-3.2-1B-Instruct and Qwen1.5-7B-Chat on individual evaluation dataset (Part A). To facilitate comparison, the best results are highlighted in bold, and the second-best results are underlined.
Model File Size(GB) SciQ(%) WinoGrande(%) ARC-E(%) ARC-C(%)
LLaMA-3.2-1B-Instruct $2.47$ $94.30$ $61.40$ $69.02$ $38.23$
[…]
Qwen1.5-7B-Chat $15.45$ $92.60$ $65.43$ $68.48$ $52.56$
[…]
TABLE X: Comparison between the proposed methods and other compressed Models of LLaMA-3.2-1B-Instruct and Qwen1.5-7B-Chat on individual evaluation dataset (Part B). To facilitate comparison, the best results are highlighted in bold, and the second-best results are underlined.
Model Hellaswag(%) BoolQ(%) NQ(%) MMLU(%) Average(%)
LLaMA-3.2-1B-Instruct $61.10$ $70.61$ $16.09$ $46.12$ $57.11$
  + SpinQuant-INT4[35] $52.17$ $62.72$ $7.87$ $37.07$ $51.18$
  + AWQ-INT4[33] $59.11$ 69.42 $12.41$ $43.36$ $55.71$
  + SmoothQuant-INT4 [52] $52.84$ $63.09$ $8.12$ $37.51$ $51.97$
  + SmoothQuant-INT8 [52] 61.08 70.52 15.73 45.90 57.02
  + Hyper-Compression 59.59 $68.78$ 14.76 43.63 55.87
Qwen1.5-7B-Chat $78.65$ $85.44$ $18.67$ $60.48$ $65.29$
  + GPTQ-INT4[20] $77.69$ $84.71$ $16.15$ 60.16 $64.12$
  + GPTQ-INT8[20] 78.57 85.44 18.64 60.52 65.23
  + Hyper-Compression 77.72 85.17 16.34 $59.24$ 64.35
[…]
No post-hoc retraining. Table XIV compares the number of tokens different compression methods use to retrain the compressed models, which highlights that our Hyper-Compression is zero-shot. This is because discrepancies between the original and decoded parameters in the Hyper-Compression are minor, at magnitudes ranging from $10^{-4}$ to $10^{-3}$. Then, the impact on error accumulation through layer-by-layer propagation is acceptable. This advantage is particularly useful in industry, where often 1) training data are inaccessible, and//or curating data is costly; 2) no computing resources are supplied for retraining. To bridge theoretical insights with practical feasibility, we empirically analyze error propagation in deep models. We conduct next-token prediction experiments using four representative models: LLaMA3.2-1B-Instruct, Qwen1.5-7B-Chat, LLaMA2-7B, and DeepSeek-R1-Distill-Qwen-14B and with four fixed input tokens (‘hyper’, ‘compression’, ‘is’, ‘magic’). For all models, we compute the Mean Absolute Error (MAE) between the layer-wise outputs of the original and compressed variants across the first 100, 200, 200, and 300 linear layers of LLaMA3.2-1B-Instruct, Qwen1.5-7B-Chat, LLaMA2-7B, and DeepSeek-R1-Distill-Qwen-14B, respectively. The empirical results provide preliminary validation of our theory, as our theory only describe the error created in a single layer.
As illustrated in Figure 10, the error accumulation surprisingly exhibits bounded fluctuation without demonstrating progressive accumulation across the network depth, indicating that our compression algorithm enjoys a stable error propagation pattern. We think this is because the trajectory of a dynamic system uniformly fills the entire space. Thus, given the weights, the compression error should also conform to an unbiased distribution, which leads to a canceling effect in Figure 10. As our current work highlights the feasibility of the proposed dynamic system-based model compression, we leave the full explanation of this phenomenon to our future work.
Figure 10: Layer-wise MAE propagation during next-token prediction on four compressed LLMs (LLaMA3.2-1B, DeepSeek-14B, LLaMA2-7B, Qwen1.5-7B) with fixed tokens (‘hyper’, ‘compression’, ‘is’, ‘magic’).
Affordable inference time. A network is a hierarchical structure. To expedite the inference time, our approach is to leverage this hierarchy by parallelizing the decoding and inference processes. The key is to complete the decoding of a layer before using this layer to infer. As shown in Table XV, “Original” refers to the time required for a single inference of the original UNet, and “Ours” refers to the time required for one inference by employing parallel decoding and inference processes. It can be seen that the Hyper-Compression only increases the total inference time moderately.
Short compression time. As Table XI shows, our method can compress models very fast. This efficiency primarily stems from using matrix operations for most compression processes and the KD-Tree to expedite the search. Additionally, by treating the compression tasks between layers as independent operations, we implement a parallel strategy to further decrease the compression time.
TABLE XI: The compression time of our method on UNet, MobileNetV3, LLaMA2-7B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B and Qwen2-72B-Instruct.
Model Time (s)
UNet + HC 30.00
MobileNetV3 + HC 11.63
Model Time (min)
LlaMA2-7B + HC 3.75
DeepSeek-R1-Distill-Qwen-14B + HC 7.79
DeepSeek-R1-Distill-Qwen-32B + HC 17.05
Qwen2-72B-Instruct + HC 33.60
6 Ablation Study and Parameter Sensitivity
As shown in Table XII, we conduct ablation experiments based on UNet and MobileNetV3 on the aforementioned two acceleration techniques to evaluate their independent contributions in computational speed. It is seen that both KD-Tree and matrix operations leads to a substantial enhancement in computational efficiency. Notably, the KD-Tree technique becomes more pronounced when the matrix operations are applied; and vice versa.
TABLE XII: The results of ablation study on acceleration techniques.
Model K-D Matrix Compression
Tree Operations Time (min)
UNet [42] 231.6
176.9
54.8
0.5
MobileNetV3 [26] 52.29
3.13
45.50
0.19
Figure 12: The experiments on synthetic weight matrices under controlled settings show that as $N_{0}$ increases, the compression error increases.
We use UNet as an example to conduct a parameter sensitivity test on three aforementioned hyperparameters: $M$, $U$, and $l$. $M$ means the maximum value of categories that is set for all layers of the model, $U$ means the list of the number of sample nodes in the center square, and $l$ means the length of the side of the center square. As shown in Table XIII, different values of $M$, $U$, and $l$ only moderately impact the model’s performance and compression ratio, which means our algorithm is robust to hyperparameters.
TABLE XIII: Parameter sensitivity results for different hyper-parameters on UNet.
Hyper-Parameters Dice ($\%$) Rate